# Vanishing criteria for Ceresa cycles

### v1

@@ 3. Vanishing criteria in the Chow and Griffiths groups
has genus $4$ and there is an isomorphism of $\mu_{3}$-representations $\operatorname{H}^{0}(C,\Omega_{C})\simeq\chi\oplus\chi\oplus\chi^{2}\oplus\chi^{2}$. It follows that $\operatorname{H}^{0}(J,\Omega_{J}^{3})^{\mu_{3}}=0$, and the abelian variety $A$ from Proposition 3.3 is 4-dimensional. If the Hodge conjecture holds for the $8$-dimensional $J\times A$, then $\bar{\kappa}(C)=0$.
@@ 4.2. Ceresa vanishing criteria
Theorem 4.2 generalizes [27, Theorem 5.16], which considered the special case where $b=0$. There, we exploited the bielliptic cover to show that the Ceresa cycle maps via a correspondence to a multiple of the point $Q_{f}=(a^{2}-4c,a(a^{2}-4c))$ on the elliptic curve $E_{f}^{\prime}\colon y^{2}=x^{3}+16\cdot{\rm disc}(f)$. This is compatible with the general case since there is a $3$-isogeny $\phi_{f}\colon E^{\prime}_{f}\to E_{f}$, and one checks using the explicit formula for $\phi_{f}$ [9, Equation (2)] that $\phi_{f}(Q_{f})=P_{f}$.
@@ 4.3. Multilinear algebra
The universal property of exterior powers induces a canonical ${\rm\mathbb{Q}}$-linear surjection $\bigwedge^{3}\operatorname{H}^{1}(J;{\rm\mathbb{Q}})\rightarrow\bigwedge^{3}_{K}\operatorname{H}^{1}(J;{\rm\mathbb{Q}})$. It is well known (see [35, Lemma 12(i)] or [17, Lemma 4.3]) that this map admits a canonical splitting, which we use to view $\bigwedge_{K}^{3}\operatorname{H}^{1}(J;{\rm\mathbb{Q}})$ as a direct summand of $\operatorname{H}^{3}(J;{\rm\mathbb{Q}})$.
@@
It suffices to prove the statements after tensoring with ${\rm\mathbb{C}}$. Let $g\in\mu_{3}\subset{\rm Aut}(C)$ be a nontrivial element. The action of $g$ on $\operatorname{H}^{1}(J;{\rm\mathbb{Q}})$ has eigenvalues (with multiplicity) $\omega,\omega,\omega,\omega^{2},\omega^{2},\omega^{2}$ so we can write $\operatorname{H}^{1}(J;{\rm\mathbb{C}})=V_{1}\oplus V_{2}$ where $V_{i}$ is the $\omega^{i}$-eigenspace. Since a three element subset of these eigenvalues have product $1$ if and only if they are all equal, $\operatorname{H}^{3}(J;{\rm\mathbb{C}})^{\mu_{3}}=(\bigwedge^{3}V_{1})\oplus(\bigwedge^{3}V_{2})$. On the other hand, an argument similar to the proof of [17, Proposition 4.4] shows $(\bigwedge^{3}_{K}\operatorname{H}^{1}(J;{\rm\mathbb{Q}}))\otimes_{K}{\rm\mathbb{C}}=\bigwedge^{3}_{K\otimes{\rm\mathbb{C}}}\operatorname{H}^{1}(J;{\rm\mathbb{C}})=(\bigwedge^{3}V_{1})\oplus(\bigwedge^{3}V_{2})$, proving the equality. The explicit description of this subspace shows that it is $2$-dimensional over ${\rm\mathbb{Q}}$. ∎
@@ 4.4. Weil classes and Schoen’s theorem
Since $J\times E$ is four-dimensional, it suffices to prove Hodge classes in $\operatorname{H}^{4}(J\times E;{\rm\mathbb{Q}})$ are algebraic. By [36, Theorem (0.1), Part (i) and (iv)], such Hodge classes are sums of products of divisor classes (which are algebraic) and Weil classes $W_{K}:=\bigwedge^{4}_{K}\operatorname{H}^{4}(J\times E;{\rm\mathbb{Q}})\subset\operatorname{H}^{4}(J\times E;{\rm\mathbb{Q}})$. Since $K={\rm\mathbb{Q}}(\omega)$ and the embedding $K\subset{\rm End}(J\times E)\otimes{\rm\mathbb{Q}}$ can be chosen to have signature $(2,2)$, Schoen has shown in [41] that the classes in $W_{K}$ are algebraic, concluding that all Hodge classes of $\operatorname{H}^{4}(J\times E;{\rm\mathbb{Q}})$ are algebraic. ∎
@@ 4.5. The Abel–Jacobi map
If $X$ is a smooth variety over ${\rm\mathbb{C}}$, we will use the notion of an integral (respectively rational) variation of (pure) Hodge structures over the complex manifold $X({\rm\mathbb{C}})$, called a ${\rm\mathbb{Z}}$-VHS (respectively ${\rm\mathbb{Q}}$-VHS) for short; see [45, §5.3.1] for definitions. If $V$ is a Hodge structure of odd weight $12k-$ 1, its intermediate Jacobian $\mathrm{J}(V)$ is the complex torus
@@
where $F^{k}\subset V\otimes_{{\rm\mathbb{Z}}}{\rm\mathbb{C}}$ is a part of the descending Hodge filtration and $V_{\tau}$ denotes the quotient of $V$ by its torsion subgroup. More generally, if $\mathbb{V}$ is a ${\rm\mathbb{Z}}$-VHS of weight $12k-$ 1 over $X({\rm\mathbb{C}})$, we can define its intermediate Jacobian $\mathrm{J}(\mathbb{V})\rightarrow X({\rm\mathbb{C}})$, a relative complex torus whose fibers over points $x\in X({\rm\mathbb{C}})$ are the classical intermediate Jacobians $\mathrm{J}(\mathbb{V}_{x})$, see [45, §7.1.1]. If $\mathbb{V},\mathbb{W}$ are two ${\rm\mathbb{Z}}$-VHS of odd weight, then $\mathrm{J}(\mathbb{V}(p))=\mathrm{J}(\mathbb{V})$ for all $p\in{\rm\mathbb{Z}}$ and a morphism $\mathbb{V}\rightarrow\mathbb{W}$ of ${\rm\mathbb{Z}}$-VHS induces a homomorphism of (relative) complex tori $\mathrm{J}(\mathbb{V})\rightarrow\mathrm{J}(\mathbb{W})$.
@@
defined in [45, §7.2.1]. Moreover, if $S/{\rm\mathbb{C}}$ is a smooth variety, $\pi\colon X\rightarrow S$ a smooth projective morphism with geometrically integral fibres and $p\in{\rm\mathbb{Z}}_{\geq 0}$, then $R^{p}\pi_{*}{\rm\mathbb{Z}}$ (pushforward of the constant sheaf in the analytic topology) has the structure of a ${\rm\mathbb{Z}}$-VHS over $S({\rm\mathbb{C}})$. If $Z$ is a codimension $p$ cycle on $X$ all of whose components are flat over $S$, and such that $Z_{s}\in{\rm CH}^{p}(X_{s})_{\hom}$ for every $s\in S({\rm\mathbb{C}})$, then Griffiths has shown that there exists a holomorphic section ${\rm AJ}(Z)$ of the relative complex torus $\mathrm{J}(R^{p}\pi_{*}{\rm\mathbb{Z}})\rightarrow S({\rm\mathbb{C}})$ with the property that ${\rm AJ}(Z)_{s}={\rm AJ}_{X}^{p}(Z_{s})$ for all $s\in S({\rm\mathbb{C}})$; this is called the normal function associated to $Z$.
@@
The first sentence follows from the compatibility of the cycle class map with correspondences [45, Proposition 9.21]. To prove the compatibility of the Abel–Jacobi map with correspondences, it suffices to prove the compatibility with pullbacks, pushforwards and intersection product. The case of pullback is elementary, using the definition of ${\rm AJ}^{p}_{X}$ in terms of extensions of Hodge structures, see [15, §2.2]. The case of pushforward follows from that of pullback and Poincare duality. Finally, compatibility with intersection product follows from [45, Proposition 9.23]. ∎
@@
Let $E$ be the elliptic curve $y^{2}=x^{3}+1$. Corollary 4.9 shows that there exists a correspondence $\gamma\in{\rm CH}^{2}(E\times J)$ such that $\gamma_{*}$ induces isomorphisms ${\rm CH}^{1}(E)_{{\rm\mathbb{Q}}}\rightarrow{\rm CH}^{2}(J)^{\mu_{3}}_{{\rm\mathbb{Q}}}$ and $\operatorname{H}^{1}(E;{\rm\mathbb{Q}})\rightarrow\operatorname{H}^{3}(J;{\rm\mathbb{Q}})^{\mu_{3}}(1)$. By Lemma 4.11 these form a commutative diagram:
@@
All arrows except the horizontal bottom one are isomorphisms of abelian groups. The same is therefore true for the bottom one ${\rm AJ}\colon{\rm CH}^{2}(J_{f})^{\mu_{3}}_{\hom,{\rm\mathbb{Q}}}\rightarrow\mathrm{J}^{2}(J_{f})\otimes{\rm\mathbb{Q}}$. Since $\kappa_{f}\in{\rm CH}^{2}(J_{f})_{\hom}^{\mu_{3}}$, we conclude that $\kappa_{f}$ is torsion if and only if ${\rm AJ}_{J_{f}}^{2}(\kappa_{f})$ is. The last claim follows from the fact that ${\rm AJ}^{2}_{J_{f}}$ is injective on torsion subgroups by a result of Murre [37, Theorem 10.3]. ∎

### v2

@@ 3. Vanishing criteria in the Chow and Griffiths groups
has genus $4$ and by [38, Lemma 2.7] we have $V(C)\simeq\chi\oplus\chi\oplus\chi^{2}\oplus\chi^{2}$ as $C_{3}$-representations. It follows that $(\bigwedge^{3}V(C))^{C_{3}}=0$ and the abelian variety $A$ from Proposition 3.3 is 4-dimensional. If the Hodge conjecture holds for the $8$-dimensional $J\times A$, then $\bar{\kappa}(C)=0$.
For our next class of examples, we consider one-parameter families of smooth projective curves with affine model of the form
$C_{t}\colon y^{m}=\left(\frac{x+1}{x-1}\right)^{a}\left(\frac{x+t}{x-t}\right)^{b},$
for integers $m,a,$ and $b$ satisfying $20<a<b<m/$ 2 and $\gcd(m,a,b)=1$. If $t\neq 0,\pm 1$, then $C_{t}$ is smooth projective of genus
$g=m+1-\gcd(a,m)-\gcd(b,m),$
which we will assume below is at least $3$. These curves admit an action by the dihedral group $D_{m}$, generated by the automorphisms $(x,y)\mapsto(x,\zeta_{m}y)$ and $\tau(x,y)=(-x,y^{-1})$.
Lemma 3.6.
The induced action of $\mu_{m}\subset D_{m}$ on $V(C_{t})$ has character $\sum_{n=1}^{m-1}\epsilon(n)\chi^{n}$, where $\chi^{n}$ is the $1$-dimensional representation of $\mu_{m}$ on which $\zeta\in\mu_{m}$ acts by $\zeta^{n}$ and
$\epsilon(n)=\begin{cases}1&\mbox{ if neither $m\nmid na$ nor $m\nmid nb$}\\ 0&\mbox{otherwise}\end{cases}$
Proof.
This follows from [38, Lemma 2.7]. ∎
Corollary 3.7.
The space $\bigwedge^{3}V(C_{t})$ has a non-trivial $D_{m}$-invariant subspace if and only if there exists a triple of integers $0<n_{1}<n_{2}<n_{3}<m$ such that for all $i\in\{1,2,3\}$ neither $m\mid n_{i}a$ nor $m\mid n_{i}b$, and such that $n_{1}+n_{2}+n_{3}=m$.
Proof.
By the lemma, a $\mu_{m}$-invariant subspace exists if and only if there exists a triple with the above divisibility properties and such that $n_{1}+n_{2}+n_{3}\in\{m,2m\}$. However, if there exists such a triple summing to $2m$ then the triple $m-n_{3},m-n_{2},m-n_{1}$ is another such triple whose sum is $m$. Finally, observe that $\tau$ swaps the $\chi^{n}$-isotypic component of $V(C_{t})$ with the $\chi^{-n}$-isotypic component. It follows that $(\wedge^{3}V(C_{t}))^{\mu_{m}}=0$ if and only if $(\wedge^{3}V(C_{t}))^{D_{m}}=0$. ∎
Below, we determine the triples $a<b<m/2$ such that $C_{t}$ satisfies the hypotheses of Theorem B.
Example 3.8.
Suppose $\gcd(m,6)=1$. As long as $m\geq 7$, the triple $31<2<m-$ 3 corresponds to an invariant subspace of $\bigwedge^{3}V(C_{t})$, so that Theorem B does not apply. However, when $m=5$, the sum of any three nonzero elements of ${\rm\mathbb{Z}}/5{\rm\mathbb{Z}}$ is nonzero and so we find a genus $4$ family which does satisfy the criteria of Theorem B, namely:
$y^{5}=\left(\frac{x+1}{x-1}\right)\left(\frac{x+t}{x-t}\right)^{2}.$
We compute that $\operatorname{H}^{3}(J)^{\mu_{5}}$ is isomorphic to $\operatorname{H}^{1}(A)(-1)$, for some abelian surface $A$ (using Proposition 3.3). By the Kani-Rosen formula [27, Theorem B], the Jacobian $J_{t}=\operatorname{Jac}(C_{t})$ is isogenous to the square of the abelian surface $\operatorname{Jac}(C_{t}/\tau)$. The Hodge conjecture is known for products of abelian surfaces [45, Theorem 3.15], so $\bar{\kappa}(C_{t})=0$ by Proposition 3.3.
Example 3.9.
If $m$ is even, then proceeding in a manner similar to the previous example, one checks that $(\wedge^{3}V(C_{t}))^{D_{m}}=0$ implies that $m=6$ (giving a subfamily of Example 3.5) or $m=12$. In the latter case, we find the family of genus $6$ curves
$y^{12}=\left(\frac{x+1}{x-1}\right)^{3}\left(\frac{x+t}{x-t}\right)^{4},$
which indeed satsifies the hypotheses of Theorem B.
Example 3.10.
In the remaining case where $m$ is an odd multiple of $3$, we check that the condition $(\wedge^{3}V(C_{t}))^{D_{m}}=0$ implies that $m=9$ or $m=15$ and moreover that one of $a$ or $b$ is $m/3$. These correspond to the genus $6$ family
(3.1) $y^{9}=\left(\frac{x+1}{x-1}\right)\left(\frac{x+t}{x-t}\right)^{3},$
and the genus $8$ family
(3.2) $y^{15}=\left(\frac{x+1}{x-1}\right)^{3}\left(\frac{x+t}{x-t}\right)^{5},$
which again satisfy the hypotheses of Theorem B.
One can check (by specializing) that the families above are generically non-hyperelliptic, so these are genuinely interesting curves satisfying $\overline{\kappa}(C)=0$ (in some cases, conditional on the Hodge conjecture). Quotienting by the involution $\tau$, we also recover the first known examples of families of curves $C$ with $\overline{\kappa}(C)=0$ and ${\rm Aut}(C)=0$:
Theorem 3.11.
Assuming the Hodge conjecture, then for $g\in\{3,4\}$, there exist non-trivial one-parameter families $\{X_{u}\}$ of genus $g$ curves over ${\rm\mathbb{C}}$ with $\bar{\kappa}(X_{u})=0$ and which generically satisfy ${\rm Aut}(X_{u})=\{1\}$.
Proof.
[…]
$X_{u}\colon y^{4}=xy^{3}+ux^{3}+3y^{2}-2ux-uy-1,$
where $u=(t-1)/(t+1)$. Specializing to $u=2$, we check in Magma that the automorphism group is trivial and hence ${\rm Aut}(X_{u})=\{1\}$ generically.
[…]
Remark 3.12.
Edgar Costa has verified for us, using [17], that the generic endomorphism algebra of $\operatorname{Jac}(X_{u})$ is contained in $K={\rm\mathbb{Q}}(\zeta_{9}+\zeta_{9}^{-1})$; by [21, §3], this inclusion is an equality. Similarly, in the genus $4$ family, the algebra ${\rm End}^{0}(\operatorname{Jac}(X_{u}))$ contains ${\rm\mathbb{Q}}(\zeta_{15}+\zeta_{15}^{-1})$.
It would be interesting to study the families in Examples 3.5 and 3.8-3.10 in more detail, and in particular to try to prove the relevant case of the Hodge conjecture and determine the locus where the Ceresa class vanishes in the Chow group (and not just the Griffiths group). We do this for the family of Picard curves, in the next section.
@@ 4.2. Ceresa vanishing criteria
Theorem 4.2 generalizes [31, Theorem 5.16], which considered the special case where $b=0$. There, we exploited the bielliptic cover to show that the Ceresa cycle maps via a correspondence to a multiple of the point $Q_{f}=(a^{2}-4c,a(a^{2}-4c))$ on the elliptic curve $E_{f}^{\prime}\colon y^{2}=x^{3}+16\cdot{\rm disc}(f)$. This is compatible with the general case since there is a $3$-isogeny $\phi_{f}\colon E^{\prime}_{f}\rightarrow E_{f}$, and one checks using the explicit formula for $\phi_{f}$ [9, Equation (2)] that $\phi_{f}(Q_{f})=P_{f}$.
@@ 4.3. Multilinear algebra
The universal property of exterior powers induces a canonical ${\rm\mathbb{Q}}$-linear surjection $\bigwedge^{3}\operatorname{H}^{1}(J;{\rm\mathbb{Q}})\rightarrow\bigwedge^{3}_{K}\operatorname{H}^{1}(J;{\rm\mathbb{Q}})$. It is well known (see [39, Lemma 12(i)] or [18, Lemma 4.3]) that this map admits a canonical splitting, which we use to view $\bigwedge_{K}^{3}\operatorname{H}^{1}(J;{\rm\mathbb{Q}})$ as a direct summand of $\operatorname{H}^{3}(J;{\rm\mathbb{Q}})$.
@@
It suffices to prove the statements after tensoring with ${\rm\mathbb{C}}$. Let $g\in\mu_{3}\subset{\rm Aut}(C)$ be a nontrivial element. The action of $g$ on $\operatorname{H}^{1}(J;{\rm\mathbb{Q}})$ has eigenvalues (with multiplicity) $\omega,\omega,\omega,\omega^{2},\omega^{2},\omega^{2}$ so we can write $\operatorname{H}^{1}(J;{\rm\mathbb{C}})=V_{1}\oplus V_{2}$ where $V_{i}$ is the $\omega^{i}$-eigenspace. Since a three element subset of these eigenvalues have product $1$ if and only if they are all equal, $\operatorname{H}^{3}(J;{\rm\mathbb{C}})^{\mu_{3}}=(\bigwedge^{3}V_{1})\oplus(\bigwedge^{3}V_{2})$. On the other hand, an argument similar to the proof of [18, Proposition 4.4] shows $(\bigwedge^{3}_{K}\operatorname{H}^{1}(J;{\rm\mathbb{Q}}))\otimes_{K}{\rm\mathbb{C}}=\bigwedge^{3}_{K\otimes{\rm\mathbb{C}}}\operatorname{H}^{1}(J;{\rm\mathbb{C}})=(\bigwedge^{3}V_{1})\oplus(\bigwedge^{3}V_{2})$, proving the equality. The explicit description of this subspace shows that it is $2$-dimensional over ${\rm\mathbb{Q}}$. ∎
@@ 4.4. Weil classes and Schoen’s theorem
Since $J\times E$ is four-dimensional, it suffices to prove Hodge classes in $\operatorname{H}^{4}(J\times E;{\rm\mathbb{Q}})$ are algebraic. By [40, Theorem (0.1), Part (i) and (iv)], such Hodge classes are sums of products of divisor classes (which are algebraic) and Weil classes $W_{K}:=\bigwedge^{4}_{K}\operatorname{H}^{4}(J\times E;{\rm\mathbb{Q}})\subset\operatorname{H}^{4}(J\times E;{\rm\mathbb{Q}})$. Since $K={\rm\mathbb{Q}}(\omega)$ and the embedding $K\subset{\rm End}(J\times E)\otimes{\rm\mathbb{Q}}$ can be chosen to have signature $(2,2)$, Schoen has shown in [46] that the classes in $W_{K}$ are algebraic, concluding that all Hodge classes of $\operatorname{H}^{4}(J\times E;{\rm\mathbb{Q}})$ are algebraic. ∎
@@ 4.5. The Abel–Jacobi map
If $X$ is a smooth variety over ${\rm\mathbb{C}}$, we will use the notion of an integral (respectively rational) variation of (pure) Hodge structures over the complex manifold $X({\rm\mathbb{C}})$, called a ${\rm\mathbb{Z}}$-VHS (respectively ${\rm\mathbb{Q}}$-VHS) for short; see [50, §5.3.1] for definitions. If $V$ is a Hodge structure of odd weight $12k-$ 1, its intermediate Jacobian $\mathrm{J}(V)$ is the complex torus
@@
where $F^{k}\subset V\otimes_{{\rm\mathbb{Z}}}{\rm\mathbb{C}}$ is a part of the descending Hodge filtration and $V_{\tau}$ denotes the quotient of $V$ by its torsion subgroup. More generally, if $\mathbb{V}$ is a ${\rm\mathbb{Z}}$-VHS of weight $12k-$ 1 over $X({\rm\mathbb{C}})$, we can define its intermediate Jacobian $\mathrm{J}(\mathbb{V})\rightarrow X({\rm\mathbb{C}})$, a relative complex torus whose fibers over points $x\in X({\rm\mathbb{C}})$ are the classical intermediate Jacobians $\mathrm{J}(\mathbb{V}_{x})$, see [50, §7.1.1]. If $\mathbb{V},\mathbb{W}$ are two ${\rm\mathbb{Z}}$-VHS of odd weight, then $\mathrm{J}(\mathbb{V}(p))=\mathrm{J}(\mathbb{V})$ for all $p\in{\rm\mathbb{Z}}$ and a morphism $\mathbb{V}\rightarrow\mathbb{W}$ of ${\rm\mathbb{Z}}$-VHS induces a homomorphism of (relative) complex tori $\mathrm{J}(\mathbb{V})\rightarrow\mathrm{J}(\mathbb{W})$.
@@
defined in [50, §7.2.1]. Moreover, if $S/{\rm\mathbb{C}}$ is a smooth variety, $\pi\colon X\rightarrow S$ a smooth projective morphism with geometrically integral fibres and $p\in{\rm\mathbb{Z}}_{\geq 0}$, then $R^{p}\pi_{*}{\rm\mathbb{Z}}$ (pushforward of the constant sheaf in the analytic topology) has the structure of a ${\rm\mathbb{Z}}$-VHS over $S({\rm\mathbb{C}})$. If $Z$ is a codimension $p$ cycle on $X$ all of whose components are flat over $S$, and such that $Z_{s}\in{\rm CH}^{p}(X_{s})_{\hom}$ for every $s\in S({\rm\mathbb{C}})$, then Griffiths has shown that there exists a holomorphic section ${\rm AJ}(Z)$ of the relative complex torus $\mathrm{J}(R^{p}\pi_{*}{\rm\mathbb{Z}})\rightarrow S({\rm\mathbb{C}})$ with the property that ${\rm AJ}(Z)_{s}={\rm AJ}_{X}^{p}(Z_{s})$ for all $s\in S({\rm\mathbb{C}})$; this is called the normal function associated to $Z$.
@@
The first sentence follows from the compatibility of the cycle class map with correspondences [50, Proposition 9.21]. To prove the compatibility of the Abel–Jacobi map with correspondences, it suffices to prove the compatibility with pullbacks, pushforwards and intersection product. The case of pullback is elementary, using the definition of ${\rm AJ}^{p}_{X}$ in terms of extensions of Hodge structures, see [15, §2.2]. The case of pushforward follows from that of pullback and Poincare duality. Finally, compatibility with intersection product follows from [50, Proposition 9.23]. ∎
@@
Let $E$ be the elliptic curve $y^{2}=x^{3}+1$. Corollary 4.9 shows that there exists a correspondence $\gamma\in{\rm CH}^{2}(E\times J)$ such that $\gamma_{*}$ induces isomorphisms ${\rm CH}^{1}(E)_{\hom,{\rm\mathbb{Q}}}\rightarrow{\rm CH}^{2}_{(1)}(J)^{\mu_{3}}$ and $\operatorname{H}^{1}(E;{\rm\mathbb{Q}})\rightarrow\operatorname{H}^{3}(J;{\rm\mathbb{Q}})^{\mu_{3}}(1)$. Let $\varphi$ be the restriction of ${\rm AJ}_{J_{f}}^{2}\otimes{\rm\mathbb{Q}}$ to ${\rm CH}_{(1)}^{2}(J_{f})^{\mu_{3}}$. By Lemma 4.11 these form a commutative diagram:
@@
All arrows except $\varphi$ are isomorphisms of abelian groups. Therefore $\varphi$ is an isomorphism too. Since the image of $\kappa_{f}$ in ${\rm CH}^{2}(J_{f})_{{\rm\mathbb{Q}}}$ lies in ${\rm CH}^{2}_{(1)}(J_{f})^{\mu_{3}}$ by (2.4), we conclude that $\kappa_{f}$ is torsion if and only if ${\rm AJ}_{J_{f}}^{2}(\kappa_{f})$ is. The claim about torsion orders follows from the fact that ${\rm AJ}^{2}_{J_{f}}$ is injective on torsion subgroups by a result of Murre [41, Theorem 10.3]. ∎